OpenStudy (anonymous):

in 35 minutes, 49 customers entered the store. At this rate, how many customers will enter the store in 2 hours?

OpenStudy (anonymous):

\[\frac{49\times 120}{35}\] and a calculator

OpenStudy (acxbox22):

since it is a constant rate you can use proportions

OpenStudy (anonymous):

how would i write it as a proportion?

OpenStudy (acxbox22):

\[\frac{ 35}{ 49 }=\frac{ 120 }{ x }\] x=number of people that enter store cross multiply and solve (might need a calculator)

OpenStudy (anonymous):

yes, that will work, but if i want to solve a proportion i would put the variable i am looking for up top in other words \[\frac{49}{35}=\frac{x}{120}\]

OpenStudy (anonymous):

\[\frac{people}{minutes}=\frac{people}{minutes}\]

OpenStudy (acxbox22):

true